40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 2 18:35:47 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
    Reply  Tue Feb 2 18:54:02 2016, Koji, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
       Reply  Wed Feb 3 23:23:47 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table KnifeEdgeScans_x.pdfBeamscan_x.pdf
          Reply  Thu Feb 4 09:16:46 2016, Koji, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
             Reply  Fri Feb 5 00:23:01 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table Beamscan_setup.pdfBeamscan_x.pdfBeamscan_y.pdfZscan.pdf
                Reply  Fri Feb 5 16:50:24 2016, gautam, Update, Green Locking, First pass at mode-matching Modematch_AUXx.pdfNewSetUp.pngModematch_AUXx_2.pdfXendModeMatch.m.zip
                   Reply  Mon Feb 8 15:36:37 2016, gautam, Update, Green Locking, Alternative mode-matching scheme Modematch_alternative.pdf
                      Reply  Tue Feb 9 04:37:10 2016, ericq, Update, Green Locking, Laser swap initiated 
                         Reply  Tue Feb 9 11:49:47 2016, gautam, Update, Green Locking, Laser swap initiated Innolight_beamProfile.pdf
                            Reply  Tue Feb 9 19:15:36 2016, gautam, Update, Green Locking, Laser swap - updates Modematch_X.pdfXendModeMatch.m.zip
                               Reply  Wed Feb 10 17:57:15 2016, gautam, Update, Green Locking, Laser swap - updates 
                                  Reply  Fri Feb 12 17:05:40 2016, gautam, Update, Green Locking, Laser swap - green light recovered but no flashes in the arm 
                                     Reply  Fri Feb 12 19:07:52 2016, Koji, Update, Green Locking, Laser swap - green light recovered but no flashes in the arm 
                                        Reply  Fri Feb 19 00:42:19 2016, gautam, Update, Green Locking, Laser swap - beam ellipticity from laser? IMG_6287.JPGIMG_6288.JPGIMG_6286.JPG
                                           Reply  Mon Feb 22 13:56:52 2016, gautam, Update, Green Locking, Laser swap -reflected beam from ETM aligned 
                                              Reply  Tue Feb 23 23:01:16 2016, gautam, Update, Green Locking, Laser swap - Green PDH locking AM_scan.pdfAM_scan_zoomed.pdf
                                                 Reply  Wed Feb 24 19:29:13 2016, gautam, Update, Green Locking, Laser swap - Green PDH locked 
                                                    Reply  Fri Feb 26 01:52:44 2016, gautam, Update, Green Locking, Laser swap - Green PDH OLTF X_PDH_OLTF_20160225.pdf
                                                       Reply  Mon Feb 29 17:17:26 2016, gautam, Update, Green Locking, Laser swap - still no green beatnote 
                                                          Reply  Wed Mar 2 17:42:19 2016, gautam, Update, Green Locking, Laser swap - some progress 
                                                             Reply  Thu Mar 3 01:25:50 2016, gautam, Update, Green Locking, Laser swap - 2 IR + 1 green beatnotes found IR_beat_20160303_2.pdf
                                                                Reply  Fri Mar 4 01:11:41 2016, gautam, Update, Green Locking, Laser swap - both green beatnotes found IR_beat_20160303_green.pdf
                                                                   Reply  Sat Mar 5 23:31:01 2016, gautam, Update, Green Locking, Laser swap - some updates AUXxTelescope.png.pngPM_setup.pdfPLLolg.pdfAMPM20160303.pdfIRbeat_20160304.pdf
                                                                      Reply  Mon Mar 7 23:51:36 2016, gautam, Update, Green Locking, Laser swap - some improvement PomonaTF.pdfXPDH.pdfYPDH.pdfgreenbeat_20160307.pdf
                                                                         Reply  Tue Mar 8 18:22:20 2016, rana, Update, Green Locking, Laser swap - some improvement 
Message ID: 12012     Entry time: Fri Feb 26 01:52:44 2016     In reply to: 12009     Reply to this: 12013
Author: gautam 
Type: Update 
Category: Green Locking 
Subject: Laser swap - Green PDH OLTF 

I spent some more time today trying to optimize the modulation frequency and amplitude for the X end PDH, and the alignment/mode-matching of the green to the arm. Some notes:

  1. After my best efforts to tweak the alignment and mode-matching into the arm by using the two lenses on translational stages, I was able to get the green TRX up to about 0.06. As mentioned in a previous elog, this is much lower than what we had with the old setup, even though we have more green power going into the arm now. However, the mode looks pretty bright and clean on the monitors. Could the large ellipticity in the beam is the limiting factor now?
  2. I measured the transfer function (attachment #1) of the PDH loop once I had settled on a modulation frequency and amplitude that I judged to be optimal (indicated on the plot). The UGF is ~7kHz. The PDH error signal as viewed on the oscilloscope is comparable to what we had with the Innolight. All this optimization was done empirically, I have yet to do the PM measurement. I can't seem to get more than 0.2 mW of IR arriving at the fiber coupler, the number I found in some older elogs is 2mW with the old setup.
  3. I did some alignment of the PSL green and the X arm green onto the beat PD on the PSL table. After the power glitches, the doubling ovens do not automatically turn on, I had turned on the end ovens earlier, and today I turned on the PSL oven. I noticed some strange behaviour initially - though the setpoint was 36.9 deg C, when I enabled the heater, there was a large overshoot (it went to almost 50deg C). I disabled the heating at this point, and re-enabled it once the oven had cooled down to ~35 deg C. I didn't observe anything like this while turning on the end ovens. But the PID parameters at the PSL table are very different, so perhaps this large overshoot and ringing is to be expected. In any case, I managed to get this working. But I was not able to find a beatnote tonight. 

To do:

  1. Verify that the two beams are aligned on the beat PD - I think I've done this carefully by checking the near and far-field, but I will double check.
  2. Find the beat note and look at the ALS noise performance with this new setup to see if it is usable even though GTRX is only 20% of what it used to be..
  3. Fix the coupling of the IR pickoff into the fiber at the endtable. Once this is done, I can do the PM measurement, and finding a beatnote may be easier given the IR beat PDs have a much wider bandwidth...
Attachment 1: X_PDH_OLTF_20160225.pdf  65 kB  | Hide | Hide all
X_PDH_OLTF_20160225.pdf
ELOG V3.1.3-